1856Dick and Barbara moved to Ispswich and purchased a home. They paid $1856 per year to insure their two cars. Dick said it cost $358 more per year to insure his car than it did to insure Barbara's car. How much didi it cost to insure Dick's car? How much did it cost to insure Barbara's car?
Answer by checkley75(3666)

x+(x+358)=1856
x+x+358=1856
2x=1856-358
2x=1498
x=1498/2
x=749 for the insurance for Barbar's car.
749+358=1107 for Dick's car.
proof
749+1107=1856
1856=1856